Harrison Company makes two products and uses a traditional costing system in which a single plantwide predetermined overhead rate is computed based on direct labor-hours. Data for the two products for the upcoming year follow: Rascon Parcel Direct materials cost per unit 14 $.7o 3.80 Dircct labor cost pcr unit Direct labor-hours per unit Number of units produced .1 IO 18.000 45.00o These products are customized to some degree for specific customers Required: 1. The company's manufacturing overhead costs for the year are expected to be $222,300. Using the company's traditional costing system, compute the unit product costs for the two products 2. Management is considering an activity-based absorption costing system in which half of the overhead would continue to be allocated based on direct labor-hours and half would be allocated based on engineering design time. This time is expected to be distributed as follows during the upcoming year: Rascon Parcel Total 4-300 4300 8,60o Enginccring design timc (in hours)
Compute the unit product costs for the two products using the proposed activity-based absorption costing system. Complete this question by entering your answers in the tabs below. Required1 Required 2 The company's manufacturing overhead costs for the year are expected to be $222,300. Using the company's traditional costing system, compute the unit product costs for the two products. (Round your intermediate calculations and final answers to 2 decimal places.) Parcel Rascon 12.57 Unit product cost $ 19.21$
Answer #1

SOLUTION

1. Total direct labor hours needed = Direct labor hours needed for Rascon + Direct labor hours needed for Parcel

= (18,000*0.10) + (45,000*0.15)

= 1,800 + 6,750 = 8,550

Total Expected manufacturing overhead costs = $222,300

Overhead rate per direct labor hour = 222,300 / 8,550 = $26

Unit product cost-

Rascon Parcel
Direct materials cost per unit 14 5
Direct labor cost per unit 2.70 3.80
Overhead cost per unit (26*0.10), (26*0.15) 2.60 3.90
Unit Product Cost 19.31 12.70

2. Calculation of Overhead allocation :

Half Overhead on the basis of traditional costing system:

Manufacturing overhead costs = (222,300/2) = 111,150

Total direct labor hours needed = 8,550

Overhead rate per direct labor hour = 111,150 / 8,550 = $13

Half Overhead on the basis of ABC system:

Manufacturing overhead costs = (222,300/2) = 111,150

Total Engineering design time = 8,600

Overhead rate per engineering hour = 111,150 / 8,600 = 12.92

Unit product cost-

Rascon Parcel
Direct materials cost per unit 14 5
Direct labor cost per unit 2.70 3.80
Overhead cost per unit:
Half Overhead on the basis of traditional costing system (13*0.10),(13*0.15) 1.30 1.95
Half Overhead on the basis of ABC system[($12.92*4,300)/18,000],[($12.92*4,300)/45,000] 3.09 1.23
Unit Product Cost 21.09 11.98